OpenStudy (anonymous):

Write 5 + 3i divided by 2 (2 + i) in the form a + bi, where a, b  R. what am i supossed to do?

OpenStudy (anonymous):

\[5+3i \div 4+2i] so then youmove the 3i to the other side by subrtacting [5\div4+i] then you move the 5 in the same way and you get -1+i\]

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

i'm not sure if that's right, so you might want to double check...

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

\[\frac{5+3i}{2(2+i)}\] multiply numerator and denominator by the conjugate of teh denominator. since the denominator is 2+i, the conjugate is 2-i so multiply numerator and denominator with that \[\implies \frac{5+3i}{2(2+i)} \times \frac{2-i}{2-i}\] \[\implies \frac{(5+3i)(2-i)}{2(2+i)(2-i)}\] simplify by FOILing \[\implies \frac{10 -5i + 6i - 3i^2}{2(4 - i^2)}\] combine like terms \[\implies \frac{10 + i - 3i^2}{2(4-i^2)}\] transform i^2 into -1 \[\implies \frac{10 + i - 3(-1)}{2(4 - (-1))}\] simplify \[\implies \frac{10 + i + 3}{2(4+1)}\] combine like terms \[\implies \frac{13+i}{2(5)}\] simplify the denominator \[\implies \frac{13+i}{10}\] did you get it?